# PEEB: Part-based Image Classifiers with an Explainable and Editable Language Bottleneck

### v1

@@ Results
| – SOTA black-box classifiers | |||
| TransFG (2022) | 86M | ViT-B/16 | 92.30 |
| ViT-Net (2022b) | 86M | DeiT-B | 93.60 |
| SR-GNN (2022) | 32M | Xception | 97.00 |
| PEEB (ours) | |||
| + trained on Stanford Dogs | 155M | OWL-ViT | 74.17 |
| + pre-trained on Dog-140 | 155M | OWL-ViT | 87.37 |
| + pre-trained on Dog-140 & | 155M | OWL-ViT | 92.20 |
| fine-tuned on Stanford Dogs | |||
@@ 6 Discussion and Conclusion
Explainability and editability PEEB stands out as a transparent and editable classifier to users by grounding the text descriptors to the visual parts in the image (Fig. 4-bottom). This transparent decision-making process plays an important role in user understanding. For instance, in Fig. 4 (upper right), the challenge of understanding why the model predicts accurately persists, particularly when we already know that the descriptors are incorrect. In contrast, PEEB’s explanations not only make errors like the mismatch between throat and wings more apparent but also enable users to adjust descriptions, thereby improving model accuracy without the need for retraining (Fig. 2).
We introduce PEEB, a pioneering part-based, explainable, and editable bird classifier that leverages textual descriptors for bird parts. By grounding natural language descriptors with visual features, PEEB brings transparency to its decision-making. Besides, PEEB achieves superior performance in both GZSL and ZSL settings compared to existing state-of-the-art explainable models. It underscores PEEB’s robustness and versatility in handling a variety of classification scenarios, especially in fine-grained tasks like avian classification. Moreover, our work contributes to the broader research community by developing the Bird-11K dataset, which encompasses a diverse range of bird species and presents a valuable resource for further explorations in fine-grained classification and beyond.
@@ 7 Limitations
Our text encoder, pre-trained on a broad image-text dataset, may not fully capture the intricate details specific to birds. Furthermore, CLIP text encoders trained by contrastive learning are known to suffer from the binding problem and do not understand some logical operators such as “and”, “or”, or negation. PEEB accuracy depends directly on the quality of the text encoder.
Dependency on image encoder for part visibility
The image encoder’s role in determining the visibility of bird parts in an image poses another limitation. Our model operates under the assumption that 12 parts are always visible in a bird image, requiring it to score these parts even when they might not be visually present. In an ideal scenario, the model should learn to assign the absence part a low score. This approach, which lacks direct supervision, relies heavily on unsupervised learning derived from class labels. Consequently, the limited dataset size of approximately 290K training images in Bird-11K may not sufficiently support robust unsupervised learning.
@@
The accuracy of our model is directly impacted by the quality of GPT-4 descriptors. Our empirical analysis across 20 bird classes revealed that, on average, 45% of these descriptors do not accurately reflect the birds’ features (Sec. F.1). However, we observed that revising certain descriptors in the CUB dataset led to a significant improvement of +10 points in classification accuracy for those classes (Sec. F.2). This primitive observation suggests that PEEB can be further improved if trained with human-labeled descriptors.
@@ Acknowledgement
We are grateful to Cornell’s Macaulay Library for providing us with 55K images categorized into 10,534 bird species for the large-scale pre-training. We thank Pooyan Rahmanzadehgervi and Giang Nguyen from Auburn University for their helpful feedback on the early results. We also thank the valuable support and feedback from Phat Nguyen. This work is supported by NaphCare Foundations, Adobe gifts, and NSF grants (1850117, 2145767). AN is supported by NaphCare Foundations, Adobe gifts, and NSF grants (1850117, 2145767).
Appendix
@@ A.2 Linear projection (for part embedding selection)
The image embedding will be forwarded to a Linear Projection layer, which is composed of a learnable logit scale, a learnable logit shift, and an Exponential Linear Unit (ELU) activation function. These processed image embeddings are then have the same dimension as the text embeddings. We select a single image embedding for each text query. In this context, the text queries correspond to the component names of the target object, which includes twelve distinct parts. This selection is based on the cosine similarity between the projected image embeddings and the text embeddings. Finally, the chosen images embeddings (before projection) will be sent to the Part MLP for classification and Box MLP for box prediction (Fig. A1, Step 1).
@@ A.3 Part MLP
We introduce Part MLP to enable part-based classification. It comprises a three-layer MLP with GELU activations (Hendrycks and Gimpel, 2016). Part MLP takes in the selected part embeddings (i.e. output of step 1 in Fig. A1) and outputs a vector of size for each part, where is the dimension of descriptor embeddings. Part MLP is trained to map the selected part embeddings to the same dimensional space with descriptor embeddings to compute final logits for classification.
@@ A.4 Box MLP
The Box MLP retained from OWL-ViT consists of a three-layer MLP. It takes the visual embedding as input and generates a four-element vector corresponding to the center coordinates and size of a bounding box (e.g., [x, y, width, height]). It is important to note that the image embedding inputs of Box MLP and Part MLP layers are the same, as shown in Fig. A1, Step 2.
@@ A.9 Computational budget and infrastructures
We use 8 Nvidia RTX A100 GPUs for our experiments. The pertaining approximate takes 24 hours on Bird-11K. The fine-tuning takes 2 to 4 hours with one single GPU.
A.10 Training objectives
As discussed in Sec. 4.3, we have three objectives during the two training stages: (a) During the pre-training stage one, we contrastively pre-train the model to maximize the similarity between related part-descriptor pairs while minimizing the unrelated pairs using symmetric cross-entropy (CE) loss (Radford et al., 2021); (b) In pre-training stage two, we try to get rid of teacher model by mimic the teacher’s choice of the boxes with symmetric cross-entropy loss; (c) In stage two, we simultaneously train PEEB to improve box prediction with DERT losses (Zheng et al., 2020).
Pre-training stage one: Symmetric cross-entropy loss for contrastive pre-training
@@ A.10 Training objectives
Pre-training stage two: Symmetric cross-entropy loss to get rid of teacher model
To get rid of the teacher model, we train PEEB to mimic the feature selection between image embedding and part name embeddings (as shown in Fig. A1, 1c). We first binary the teacher logits and consider it as the ground truth label. Then, apply the same symmetric cross-entropy loss as described in eq. 7 with two minor differences: (1) The text input is part names rather than descriptions. (2) The Part MLP is replaced by Linear Projection (Fig. A1, 2c).
Fine-tuning: DETR losses
@@
Step 2: (a) Encode input texts (i.e. N sets of 12-part descriptors) with the same text encoder to get . (b) Feed the selected part embeddings to box MLP to localize parts (in center format). (c) Also feed the selected part embeddings to part MLP to get in the same dimensional space with (d) Compute dot product between and , then diagonal sum for each class and over logits to get predicted label .
@@ B.2 Model notations
We named fine-tuned models after the pre-trained model and the fine-tuned training set. For example, PEEB is fine-tuned from PEEB, on CUB training set.
@@ Appendix C Generating part-based descriptors
Where {class name} is substituted for a given bird name.
@@ Appendix D Datasets
D.1 Bird-11K
We provide a brief statistic of Bird-11K in Table A4. Bird-11K is a diverse and long-tailed avian dataset that only includes bird images. The descriptors generated by GTP4 are in English and only describe the visual features of the corresponding class. We propose Bird-11K for academic research only.
| Dataset | Images | Species |
@@ D.1 Bird-11K
| iNaturalist 2021-birds (Van Horn et al., 2021) | 74,300 | 1,464 |
@@
NABirds v1 (Van Horn et al., 2015): For non-commercial research purposes, other use is restricted 333See Terms of Use here for detail: .
Birdsnap v7 (Berg et al., 2014): The dataset creator provides no specific license or terms of use. We only use this dataset for academic research until more specific details can be obtained.
iNaturalist 2021-birds (Van Horn et al., 2021): CC0: Public Domain.
@@
D.2 Dog-140
To pre-train PEEB on dogs, we construct Dog-140 by combining dog images from ImageNet21K and Stanford Dogs. Specifically, we selected 189 dog classes from ImageNet21K, and based on Fédération Cynologique Internationale (FCI) (Fédération Cynologique Internationale (2023), FCI), we merged them with 120 classes from Stanford Dogs, ending up with 140 classes. After merging, Dog-140 has 206,076 images in total. We provide a class distribution analysis in Fig. A4, where we can find that Dog-140 is roughly class-balanced.
@@ Results
The primary objective is to ascertain the superiority of our pre-trained model, PEEB, against benchmarks like CLIP and M&V. For CUB, our method reported a classification accuracy of 17.9%, contrasting the 5.95% and 7.66% achieved by CLIP and M&V, respectively, as shown in Table A5. The PEEB score, which is marginally higher (+10) than M&V, highlights the advantages of our method that utilizes component-based classification. On the NABirds, our method surpassed the CLIP and M&V by (+1) point. The performance disparity between CUB and NABirds can be attributed to two factors: the elevated complexity of the task (555-way classification for NABirds versus 200-way for CUB) and the marked reduction in training data. An auxiliary observation, detailed in Sec. E.3, indicates that our pre-trained model necessitates at least 250k images to achieve admirable classification accuracy on CUB, but we only have 210k images training images in Bird-11K (Table A3).
@@ E.7 Evaluation of predicted boxes from PEEB
Given our focus on part-based classification, we aimed to ascertain the quality of our model’s box predictions. To this end, we employed two metrics: mean Intersection over Union (IoU) and precision based on key points. We opted for mean IoU over the conventional mAP because: (1) Ground-truth boxes for bird parts are absent, and (2) our model is constrained to predict a single box per part, ensuring a recall of one. Thus, we treat OWL-ViT’s boxes as the ground truth and evaluate the box overlap through mean IoU. Furthermore, leveraging human-annotated key points for bird parts, we measure the precision of predicted boxes by determining if they contain the corresponding key points. We evaluate our fine-tuned models on their corresponding test sets. For instance, PEEB, fine-tuned based on the CUB split (Akata et al., 2015), is evaluated on the CUB test set.
@@ Appendix F Study on GPT-4 generated descriptors
F.1 Noise measurement in GPT-4 generated descriptors
@@
F.2 Revising descriptors improves classification accuracy
@@
Due to the errors in the descriptors we used to train PEEB, simply replacing the descriptors with their revised version does not lead to better performance. Because the incorrect descriptors in training change the meaning of some of the phrases. For example, the belly of Blue bunting is pure blue, but the descriptors from GPT-4 is soft, creamy white. In addition, the GTP-4 uses the exact same descriptor in the belly for other classes, e.g., Blue breasted quail, which should be cinnamon. Blue Fronted Flycatcher, which should be yellow. Training the same descriptors with different colors confuses the model, and the model will convey the phrase “creamy white” with a different meaning to humans. Therefore, simply changing the descriptors to their’ revised version will not work. We empirically inspect the descriptors that PEEB can correctly respond to and replace the class descriptors with the revised version. Specifically, we replace the descriptors of 17 classes in CUB and test the classification accuracy on PEEB.
@@ Results
As shown in Table A12, the overall accuracy increase +0.8 points. The average improvement of the revised class is around +10.8, hitting that if we have correct descriptors of all classes, we may significantly improve the classification accuracy of the pre-trained model. However, correcting all 11k class descriptors is too expensive and out of the scope of this work. We leave it as a further direction of improving the part-based bird classification.
@@ G.4 Examples of PEEB explanations for dogs
Figs. A13, A14, A15, A16, A17, A18 and A19 are examples of how PEEB makes classification based on the descriptors in Stanford Dogs dataset. We showed that our model works well on dogs, which indicates that our proposed method is transferable to other domains while maintaining nice explainability as in birds.

### v3

@@ 5.6 Applying PEEB to
| PEEB (ours) | 155M | ||
| pre-training + finetuning only | 155M | OWL-ViT | 74.17 |
| pre-training + finetuning | 155M | OWL-ViT | 87.37 |
| pre-training + finetuning | 155M | OWL-ViT | 92.20 |
@@ 6 Discussion and Conclusion
We introduce PEEB, a unique, novel explainable classifier due to its editability (Fig. 2) and operation at the part level on both image and text sides. The part-level operation makes PEEB applicable to fine-grained classification. Yet, it is also interesting to extend PEEB into an object-level model for multi-domain tasks like ImageNet or VQA.
Besides enabling users to edit PEEB’s text descriptors to re-program PEEB, it might also be promising to let users edit the bounding boxes while working with PEEB to improve the human-AI team accuracy Nguyen et al. (2024). On object detection, PEEB’s Box MLP performs on-par with OWL-ViT based on quantitative (Sec. E.7) and qualitative results (Appendix G).
Finally, we contribute to the broader research community by curating the Bird-11K and Dog-140 datasets and showing that it is possible to leverage them for large-scale training.
@@ 7 Limitations
Our CLIP text encoder, pre-trained on an Internet-scale image-text dataset Radford et al. (2021), may not fully capture the intricate details specific to birds. Furthermore, the CLIP text encoder is known to suffer from the binding problem and do not understand some logical operators such as “and”, “or”, or negation. PEEB accuracy depends heavily on the quality of the text encoder.
Assumption that object parts mostly visible
PEEB operates based on the assumption that most if not all of the object parts are visible in the image. In cases where a part is missing or occluded, the model may still assign a non-zero similarity score (i.e. a non-zero dot product between the image-part embedding and its associated text descriptor), which makes it harder to separate classes. It might be beneficial to incorporate extra training samples and specifically encourages PEEB to assign zero image-text similarity score to the missing or occluded parts.
@@
The accuracy of PEEB is directly governed by the accuracy of descriptors, which are currently generated by GPT-4. Yet, our manual assessment over 20 bird classes reveals that, on average, 45% of these descriptors do not accurately reflect the birds’ features (Sec. F.2). Also, we observe that revising certain descriptors in the CUB dataset led to a significant improvement of +10 points in classification accuracy for those classes (Sec. F.3). This primitive observation suggests that PEEB can be further improved if trained with more accurate, human-labeled descriptors.
@@ Acknowledgement
We are grateful to Cornell’s Macaulay Library for providing us with 55K images categorized into 10,534 bird species for the large-scale pre-training. We thank Pooyan Rahmanzadehgervi, Giang Nguyen, Tin Nguyen, and Hung Huy Nguyen from Auburn University for their helpful feedback on our early results. We also thank Phat Nguyen for her valuable support and feedback. AN is supported by NaphCare Foundations, Adobe gifts, and NSF grant no. 2145767.
Appendix for:
PEEB: Part-based Image Classifiers with an
Explainable and Editable Language Bottleneck
@@ A.2 Linear projection (for part embedding selection)
The image embedding will be forwarded to a Linear Projection layer (see detail implementation here), which is composed of a learnable logit scale, a learnable logit shift, and an Exponential Linear Unit (ELU) activation function. These processed image embeddings then have the same dimension as the text embeddings. For OWL-ViT, the image embeddings are projected from 768 to 512. We select a single image embedding for each text query. In this context, the text queries correspond to the component names of the target object, which includes twelve distinct parts. This selection is based on the cosine similarity between the projected image embeddings and the text embeddings. Finally, the chosen images embeddings (before projection) will be sent to the Part MLP for classification and Box MLP for box prediction (Fig. A1, Step 1).
@@ A.3 Part MLP
We introduce Part MLP to enable part-based classification (see implementation detail here). It comprises a three-layer MLP with GELU activations (Hendrycks and Gimpel, 2016) . Part MLP takes in the selected part embeddings (i.e. output of step 1 in Fig. A1) and outputs a vector of size for each part, where is the dimension of descriptor embeddings (for OWL-ViT, the input dimension is 768, and ). Part MLP is trained to map the selected part embeddings to the same dimensional space with descriptor embeddings to compute final logits for classification.
@@ A.4 Box MLP
The Box MLP retained from OWL-ViT consists of a three-layer MLP (see here for implementation detail). It takes the visual embedding as input and generates a four-element vector corresponding to the center coordinates and size of a bounding box (e.g., [x, y, width, height]). It is important to note that the image embedding inputs of Box MLP and Part MLP layers are the same, as shown in Fig. A1, Step 2.
@@ A.5 Visual part embedding selection
Step 1: (a) Encode an input image and texts (i.e. 12 part names) by the image and text encoder to get patch embeddings and text embeddings . (b) Feed to Linear Projection to get in the same dimensional space with and compute dot product between and . (c) over embeddings to select 12 part embeddings.
Step 2: (a) Encode input texts (i.e. N sets of 12-part descriptors) with the same text encoder to get . (b) Feed the selected part embeddings to Box MLP to localize parts (in center format). (c) Also feed the selected part embeddings to Part MLP to get in the same dimensional space with (d) Compute a dot product between and , then diagonal sum for each class and over logits to get predicted label .
@@ A.9 Computational budget and infrastructures
We use 8 Nvidia RTX A100 GPUs for our experiments. The pertaining approximate takes 24 hours on Bird-11K. The finetuning takes 2 to 4 hours with one single GPU.
A.10 Pre-training and Finetuning objectives
As discussed in Sec. 4.3, we have three objectives during the Pre-training phase:
- 1.
2.
Pre-training Stage 2: (Fig. A3) We try to remove the dependence on the OWL-ViT teacher model by training PEEB to mimic OWL-ViT’s box predictions using the SCE loss.
- 3.
During the Finetuning phase where we finetune on a downstream task (e.g. Dogs-120 or CUB-200), we also employ the same three losses. However, we change the first loss from SCE into CE since on the downstream classification task, the classifier is tasked with selecting one class that matches the single input image from a set of classes.
A.10.1 Pre-training stage one: Symmetric cross-entropy loss for contrastive pre-training
@@
A.10.2 Pre-training stage 2: Symmetric cross-entropy loss to mimic the teacher OWL-ViT detector
To mimic the object detection capability of the OWL-ViT teacher, we train PEEB to mimic the image-text similarity prediction between image embedding and textual part-name embeddings (as shown in Fig. A1, 1c). We first binary the teacher logits and consider it as the ground truth label. Then, apply the same symmetric cross-entropy loss as described in eq. 7 with two minor differences: (1) The text input is part names rather than descriptions. (2) The Part MLP is replaced by Linear Projection (Fig. A1, 2c).
A.10.3 Pre-training stage 2: DETR losses to mimic the teacher OWL-ViT detector
@@ B.2 Model notations
We named finetuned models after the pre-trained model and the finetuned training set. For example, PEEB is finetuned from PEEB, on CUB training set.
@@ Appendix C Generating part-based descriptors
Where {class name} is substituted for a given bird name (e.g., Painted Bunting).
@@ Appendix D Datasets
D.1
Bird-11K
We provide a brief statistic of Bird-11K in Table A4. Bird-11K is a diverse and long-tailed bird-image dataset. The descriptors generated by GPT-4 are in English and only describe the visual features of the corresponding class. We propose Bird-11K for academic research only.
| Dataset | # of Images | # of Species |
@@
| iNaturalist 2021-birds (Van Horn et al., 2021) | 74,300 | 1,320 |
@@
NABirds-v1 (Van Horn et al., 2015): For non-commercial research purposes, other use is restricted 333See Terms of Use here for detail: .
Birdsnap-v7 (Berg et al., 2014): The dataset creator provides no specific license or terms of use. We only use this dataset for academic research until more specific details can be obtained.
@@
D.2
Dog-140
To pre-train PEEB on dogs, we construct Dog-140 by combining dog images from ImageNet-21K and Stanford Dogs-120. Specifically, we selected 189 dog classes from ImageNet-21K, and based on Fédération Cynologique Internationale (FCI) (Fédération Cynologique Internationale (2023), FCI), we merged them with 120 classes from Stanford Dogs, ending up with 140 classes. After merging, Dog-140 has 206,076 images in total. We provide a class distribution analysis in Fig. A4, where we can find that Dog-140 is roughly class-balanced.
@@ Results
The primary objective is to ascertain the superiority of our pre-trained model, PEEB, against benchmarks like CLIP and M&V. For CUB, our method reported a classification accuracy of 17.9%, contrasting the 5.95% and 7.66% achieved by CLIP and M&V, respectively, as shown in Table A5. The PEEB score, which is substantially higher (+10) than M&V, highlights the advantages of our part-based classification. On NABirds, our method surpasses CLIP and M&V by +1 point. The performance disparity between CUB and NABirds can be attributed to two factors: The elevated complexity of the task (555-way classification for NABirds versus 200-way for CUB) and the marked reduction in training data. An auxiliary observation, detailed in Sec. E.3, indicates that our pre-trained model necessitates at least 250k images to achieve admirable classification accuracy on CUB, but we only have 210k images training images in Bird-11K (the variants of Bird-11K with classes excluded for ZSL testing are described in Table A3).
@@ E.7 Evaluation of predicted boxes from PEEB
Given our focus on part-based classification, we aimed to ascertain the quality of our model’s box predictions. To this end, we employed two metrics: mean Intersection over Union (IoU) and precision based on key points. We opted for mean IoU over the conventional mAP because: (1) Ground-truth boxes for bird parts are absent, and (2) our model is constrained to predict a single box per part, ensuring a recall of one. Thus, we treat OWL-ViT’s boxes as the ground truth and evaluate the box overlap through mean IoU. Furthermore, leveraging human-annotated key points for bird parts, we measure the precision of predicted boxes by determining if they contain the corresponding key points. We evaluate our finetuned models on their corresponding test sets. For instance, PEEB, finetuned based on the CUB split (Akata et al., 2015), is evaluated on the CUB test set.
@@ Appendix F Study on GPT-4 generated descriptors
F.1 Assessment of the generated part-based descriptors
We test GPT-4V on the CUB test set using the generated descriptors of 200 classes to assess their usability. Specifically, we feed GPT-4V with each test image encoded in the payload and 200 sets of part-based descriptors through a carefully designed prompt (Table A11). GPT-4V is asked to output one of 200 provided class names to compute the classification accuracy. As a result, GPT-4V achieves 69.4% accuracy which is slightly higher than PEEB’s generalized zero-shot accuracy (64.33%) and significantly lower than PEEB results after finetuning (86-88%).
F.2 Noise measurement in GPT-4 generated descriptors
@@
F.3 Revising descriptors improves classification accuracy
@@
Due to the errors in the descriptors we used to train PEEB, simply replacing the descriptors with their revised version does not lead to better performance. Because the incorrect descriptors in training change the meaning of some of the phrases. For example, the belly of Blue bunting is pure blue, but the descriptors from GPT-4 is soft, creamy white. In addition, the GPT-4 uses the exact same descriptor in the belly for other classes, e.g., Blue breasted quail, which should be cinnamon. Blue Fronted Flycatcher, which should be yellow. Training the same descriptors with different colors confuses the model, and the model will convey the phrase “creamy white” with a different meaning to humans. Therefore, simply changing the descriptors to their’ revised version will not work. We empirically inspect the descriptors that PEEB can correctly respond to and replace the class descriptors with the revised version. Specifically, we replace the descriptors of 17 classes in CUB and test the classification accuracy on PEEB.
@@ Results
As shown in Table A13, the overall accuracy increases by +0.8 points.
The average improvement of the revised class is around +10.8, hitting that if we have correct descriptors of all classes, we may significantly improve the classification accuracy of the pre-trained model. However, correcting all 11k class descriptors is too expensive and out of the scope of this work. We leave it as a further direction of improving the part-based bird classification.
@@ G.4 Examples of PEEB explanations for dogs
Figs. A13, A14 and A15 are examples of how PEEB makes classification based on the descriptors in Stanford Dogs dataset. We demonstrate that our model works well on dogs, which indicates that our proposed method is transferable to other domains while maintaining high-quality explainability as in birds.